find the slope of the line that contains the point (-1,5) and (5,-4).

Answer by Alan3354(69443)

find the slope of the line that contains the point (-1,5) and (5,-4).
---------
Slope, m = diffy/diffx
m = (5 - (-4))/(-1-5) = 9/-6
m = -3/2 or -1.5